Pina Co. purchased for $2,422,000 property that incuded both land and a building to be used in operations. The sellers book value was $304,000 for the land and $855,000 for the building. By appraisal, the fair value was estimated to be $818,670 for the land and $1,910,230 for the building. At what amount should Pina report the land and the building at the end of the year? (Round intermediate calculations to 5 decimal places, e.g. 1.25124 and final answers to o decimal places e.g. 58,971.) Cost allocated to land Cost allocated to building
Answer #1

cost should be allocated to land and building on the basis of fair value of land and building:

here, the fair value to be allocated = (818,670 + 1,910,230) =>2,728,900

value to be allocated to land = 2,422,000 * (1,910,230 / 2,728,900)

=>1,695,400

value to be allocated to building = 2,422,000 * (818,670 / 2,728,900)

=>726,600.

cost allocated to land 1,695,400
cost allocated to building 726,600